# From multitype branching Brownian motions to branching Markov additive processes

### v1

@@ 1.1 A motivating model: on-off branching Brownian motions
The starting point of this work is a model introduced by Blath et al. in [BlaEtAl25]. Motivated by the biological concepts of dormancy and seed banks the authors proposed a branching particle system with two types: active and dormant. The two types are distinguished by different branching rates, reproduction laws, and motion dynamics. Furthermore, particles switch between the active and dormant states at certain rates. This system can be viewed as an extension of the well-known (single-type) branching Brownian motion (BBM) and is called on-off BBMs. Similar to the classical BBM, the on-off BBMs are connected to travelling wave solutions of two-type Fisher-Kolmogorov-Petrovskii-Piskounov (FKPP) equations, featured with dormancy mechanisms. While [BlaEtAl25] establishes the convergence of additive martingales and the existence of travelling waves in the supercritical regime, several problems remain open, including the following ones posed by these authors [BlaEtAl25, Section 4]:
  • The uniqueness of the travelling waves.
  • A probabilistic representation of the travelling waves akin to the Lalley-Sellke [LalleySellke87] construction.
@@
A classical technique for addressing such questions is the spine decomposition, which is a seminal method in branching process theory (see e.g. [ShiBRW]). However, its application to the on-off BBMs is non-trivial. Although the literature on single-type and multitype BBMs is vast, it usually assumes the underlying particle motion is identical for all types; this is crucially different from the on/off BBMs, where the motions of particles are type-dependent. This seemingly minor extension introduces non-trivial technical difficulties. As Blath et al. note [BlaEtAl25], “the quadratic variation is truly probabilistic, making an application of the Girsanov Theorem difficult.” This inherent difficulty has thus posed a major obstacle to adapting the powerful spine decomposition technique.
One initial motivation for this work is to overcome this obstacle. We find that the essence of the problem is easier captured when we shift to a more general framework: branching particle systems with type-dependent Lévy processes. This perspective is naturally related to Markov additive processes (MAPs), which provide a clearer view of the underlying structure and the necessary mathematical tools to handle the heterogeneity. In this framework, we introduce a new spine decomposition. This allows us to provide a systematic approach to studying such branching systems and, in particular, resolve several open problems identified by Blath et al. [BlaEtAl25] mentioned above.
@@ 1.2 Multitype branching Lévy processes viewed as branching MAPs
Our models are multitype branching Lévy processes with a finite type space $\mathcal{I}=\{1,\ldots,\mathtt{d}\}$, for some $\mathtt{d}\in\mathbb{N}$. Each particle is assigned a type $i\in\mathcal{I}$ and moves in $\mathbb{R}$ according to a type-dependent Lévy process. They further randomly either branches, generating offspring that all inherit its current type, or switches to a new type. All particles evolve independently of one another. Specifically, the model is described as follows.
  • (Movement) For each particle of type $i\in\mathcal{I}$, its movement is governed by a $\mathbb{R}$-valued Lévy process $(\chi_{i}(t),t\geq 0)$. For $\theta\geq 0$, when the exponential moment $\mathbb{E}[e^{-\theta\chi_{i}(1)}]$ is finite, we define the Laplace exponent $\phi_{i}$ such that
@@
    $\mathcal{A}_{i}f(x)=\frac{1}{2}\sigma_{i}^{2}\partial_{xx}f(x)+\mathtt{a}_{i}\partial_{x}f(x)+\int_{\mathbb{R}}\Big[f(x+y)-f(x)-y\mathbbm{1}_{\left\{|y|<1\right\}}\partial_{x}f(x)\Big]\,\Lambda_{i}(\mathrm{d}y).$ (1.3)
@@
  • (Switching types) Each particle changes its type according to a continuous-time Markov chain $\Theta$ on $\mathcal{I}$, with intensity matrix $Q=(q_{ij})_{1\leq i,j\leq N}$, and possibly makes a jump at the time when it changes type. More precisely, at rate $q_{ij}\geq 0$, a particle of type $i$ switches to the type $j\in\mathcal{I}$, and at the same time makes a jump in space according to the law of a real-valued random variable $U_{ij}$, with convention that $U_{ij}=0$ if $q_{ij}=0$, and that $U_{ii}=0$. For $i\in\mathcal{I}$, let $q_{i}=-q_{ii}=\sum_{j\neq i}q_{ij}$. When $\mathbb{E}[e^{-\theta U_{ij}}]$ is finite, we define
@@
The theory of MAPs is well-established and has prominent applications in e.g. classical applied probabilistic models for queues. We refer to [Asmussen, Chapter XI], [Iva2011, Chapter 2] and [DLK17, Appendix] for detailed discussions. The following proposition is well-known, giving a standard representation for MAPs.
@@ 1.3 Main results
Let $\theta\in[0,\bar{\theta})$. By the Perron-Frobenius (PF) theorem (see [seneta06, Theorem 1.1] or [Horn_Johnson, Theorem 8.3.4]), since $M^{(\theta)}(t)$ is a matrix with positive entries, the matrix $\mathcal{M}(\theta)$ has a PF eigenvalue $\lambda(\theta)$. This eigenvalue $\lambda(\theta)$ is real and larger than the real part of any other eigenvalues of $\mathcal{M}(\theta)$, and its associated right eigenvector $\vec{V}(\theta)=(V_{i}(\theta),i\in\mathcal{I})$ has strictly positive entries. Furthermore, because the entries of $\mathcal{M}(\theta)$ are infinitely differentiable on $(0,\bar{\theta})$, the PF eigenvalue $\lambda(\theta)$ and the corresponding eigenvector $\vec{V}(\theta)$ are also infinitely differentiable on this interval. The properties above of the PF eigenvalue are introduced in Lemma 2.3 and Lemma 2.4.
We observe that the number of particles $\left(\sum_{u\in\mathcal{N}_{t}}\mathbbm{1}_{\left\{J_{u}(t)=1\right\}},\ldots\sum_{u\in\mathcal{N}_{t}}\mathbbm{1}_{\left\{J_{u}(t)=\mathtt{d}\right\}}\right)_{t\geq 0}$, forms a continuous-time multitype branching process (see e.g. [Athreya1968]): for each particle of type $i\in\mathcal{I}$, it is replaced by $k$ particles of type $i$ at rate $\beta_{i}\mu_{i}(k)$ for $k\geq 0$; whereas it is replaced by one type $j$ particle at rate $q_{ij}$ for $j\neq i$. When $\theta=0$, $\mathcal{M}(0)=Q+\mathrm{diag}(\beta_{1}m_{1}-\beta_{1},\ldots,\beta_{\mathtt{d}}m_{\mathtt{d}}-\beta_{\mathtt{d}})$ serves as the generator of the first moment semigroup of $\left(\sum_{u\in\mathcal{N}_{t}}\mathbbm{1}_{\left\{J_{u}(t)=1\right\}},\ldots\sum_{u\in\mathcal{N}_{t}}\mathbbm{1}_{\left\{J_{u}(t)=\mathtt{d}\right\}}\right)_{t\geq 0}$ and the PF eigenvalue $\lambda(0)$ determines the extinction behaviour with a phase transition. Denote the survival event by
@@
By [Athreya1968, Theorem 2], the vector $\vec{\mathbf{q}}:=(\mathbf{q}_{1},\dots,\mathbf{q}_{\mathtt{d}})$ is the unique solution in $0,1]^{\mathtt{d}}\setminus\{(1,\dots,1)\}$ to the equation:
@@
The process also satisfies the “positive regularity” and “non-singularity” in [Harris63, Theorem 2.7.1] and [Harris63, Corollary 1 of Theorem 2.7.2]. Therefore the unique solution $\vec{\mathbf{q}}$ also lies in $0,1)^{\mathtt{d}}$. Whereas when $\lambda(0)\leq 0$, the extinction happens $\mathbf{P}_{x,i}$-a.s. $\forall x\in\mathbb{R},i\in\mathcal{I}$, i.e. $\mathbf{P}_{x,i}(\mathscr{S})=0$.
In what follows, we will always work in the supercritical regime $\lambda(0)>0$. Summarizing, we assume that
@@
$W_{\theta}(t):=\sum_{u\in\mathcal{N}_{t}}e^{-\left(\theta X_{u}(t)+\lambda(\theta)t\right)}V_{J_{u}(t)}(\theta).$ (1.12)
@@
Theorem 1.4 ($\mathcal{L}^{1}$-convergence).
@@
By the main Theorem of [Kingman61], the function $q\mapsto\lambda(q)$ is strictly convex on $(0,\bar{\theta})$. It follows that $q\mapsto\frac{\lambda(q)}{q}$ either attains its unique minimum in the interval $(0,\bar{\theta})$ or approaches the infimum at the boundaries 0 or $\bar{\theta}$ (note that $\bar{\theta}$ could be $\infty$). The next assumption specifies that the minimum is attained in the interior.
@@
The main tool we use in the proof of Theorem 1.4 is the spine decomposition. The spine decomposition is a powerful technique in the study of branching processes. Typically, it contains two steps. The first step is a change of measure, and the second step is a deconstruction of the process in the new probability measure. For the details of the spine decomposition on classic branching systems, one can refer to [Lyons97, ShiBRW] for applications on branching random walks (BRWs), [KYPRIANOU04] for applications on branching Brownian motions (BBMs), and [RS20] for applications on branching Markov processes (BMPs). In our model, we established a spine decomposition, which is introduced formally in Section 2.3, and proved in Section 5. This framework provides a powerful tool for deeper analysis of the model, which we employ here to study the derivative martingale and the travelling waves.
@@
$Z_{\theta}(t):=\sum_{u\in\mathcal{N}_{t}}e^{-\theta X_{u}(t)-\lambda(\theta)t}\left[V_{J_{u}(t)}(\theta)(X_{u}(t)+\lambda^{\prime}(\theta)t)-V’_{J_{u}(t)}(\theta)\right],\qquad t\geq 0.$ (1.13)
Note that we intuitively have $Z_{\theta}(t):=-\partial_{\theta}W_{\theta}(t)$. At criticality $\theta=\theta^{*}$, the derivative martingale is a crucial object for understanding the fine structure of the frontier in branching systems; see, for example, [Aidekon13, ShiBRW] for applications on branching random walks, [KYPRIANOU04, LalleySellke87] for applications on BBMs and [Ren2014] for applications on multitype BBMs.
@@
The FKPP equation is a well-known reaction-diffusion equation that appears in population genetics and ecology. Its connection to branching processes was first studied via probabilistic tools in [Mckean75]. Since then, it has been widely studied in probability, particularly in branching systems, see, for example, [Champneys95], [KLMR12], [KYPRIANOU04], [LalleySellke87] and [Ren2014]. In the study of FKPP equation, the existence, uniqueness and the asymptotic of the travelling wave solutions are particularly interested. One can find a “common pattern” in these works of FKPP equations. In the supercritical regime, the travelling wave solutions are often strongly connected to the limit of the additive martingale; while in the critical regime, it is related to the limit of the derivative martingale; while in the subcritical regime, there is no travelling wave solution.
@@
$\frac{\partial\mathbf{u}(t,x,i)}{\partial t}=\mathcal{A}_{i}\mathbf{u}(t,x,i)+\sum_{j\neq i}q_{ij}\left(\int_{\mathbb{R}}\mathbf{u}(t,x\!+\!y,j)\mathbb{P}(U_{ij}\!\in\!\mathrm{d}y)-\mathbf{u}(t,x,i)\right)+\beta_{i}\big(g_{i}(\mathbf{u}(t,x,i))-\mathbf{u}(t,x,i)\big).$ (1.15)
@@
$\mathbf{u}(t,x,i)=\mathbb{E}_{x,i}\left[\mathbf{u}(0,\xi_{i}(t),i)\right]+\int_{0}^{t}\mathbb{E}_{x,i}\Bigg[$ $\beta_{i}g_{i}\left(\mathbf{u}\left(s,\xi_{i}(t\!-\!s),i\right)\right)+\sum_{j\neq i}q_{ij}\int_{\mathbb{R}}\mathbf{u}\left(s,y\!+\!\xi_{i}(t\!-\!s),j\right)\mathbb{P}(U_{ij}\in\mathrm{d}y)$
$-(q_{i}+\beta_{i})\mathbf{u}\left(s,\xi_{i}(t\!-\!s),i\right)\Bigg]\mathrm{d}s.$ (1.16)
@@
$\mathcal{T}_{1}=\left\{f\;\middle|\;\forall i\in\mathcal{I},x\mapsto f(x,i)\text{ is non-decreasing in }x,\lim_{x\to-\infty}f(x,i)=\mathbf{q}_{i}\text{ and }\lim_{x\to\infty}f(x,i)=1\right\}.$ (1.17)
@@
    (Subcritical regime) If $\rho<\frac{\lambda(\theta^{*})}{\theta^{*}}$, there are no travelling wave solutions with speed $\rho$.

### v2

@@ 1.1 A motivating model: on-off branching Brownian motions
The starting point of this work is an elegant two-type branching Brownian motion model introduced by Blath et al. [10]. Their model, so-called on-off branching Brownian motion (on-off BBM), incorporates the biological concepts of dormancy and seed banks. In this branching system, particles exist in either an active or dormant state, with each type following distinct branching rates, reproduction laws, and motion dynamics. The model also allows stochastic switching between the two states. This framework provides a natural extension of the well-known (single-type) branching Brownian motion. As in the classical case, an on-off BBM is intimately related to a two-type extension of Fisher–Kolmogorov–Petrovskii–Piskunov (FKPP) equation with the dormancy feature. Through a careful and insightful analysis, [10] established two key results in the supercritical regime: the convergence of additive martingales and the existence of travelling waves. Moreover, the authors highlighted several open questions in [10, Section 4], including the following:
  • The uniqueness of the travelling waves;
  • A probabilistic representation of the travelling waves akin to the Lalley-Sellke [38] construction;
@@
A classical technique for addressing such questions is the spine decomposition, which is a seminal method in branching process theory (see e.g. [54]). However, its application to the on-off BBMs is non-trivial. Although the literature on single-type and multitype BBMs is vast, it usually assumes the underlying particle motion is identical for all types; this is crucially different from the on/off BBMs, where the motions of particles are type-dependent. This seemingly minor extension introduces non-trivial technical difficulties. As Blath et al. note [10], “the quadratic variation is truly probabilistic, making an application of the Girsanov Theorem difficult.” This inherent difficulty has thus posed a major obstacle to adapting the powerful spine decomposition technique.
One initial motivation for this work is to overcome this obstacle. We find that the essence of the problem is easier captured when we shift to a more general framework: branching particle systems with type-dependent Lévy processes. This perspective is naturally related to Markov additive processes (MAPs), which provide a clearer view of the underlying structure and the necessary mathematical tools to handle the heterogeneity. In this framework, we introduce a new spine decomposition. This allows us to provide a systematic approach to studying such branching systems and, in particular, resolve several open problems identified by Blath et al. [10] mentioned above.
@@ 1.2 Multitype branching Lévy processes viewed as branching MAPs
Our models are multitype branching Lévy processes with a finite type space $\mathcal{I}=\{1,\ldots,\mathtt{d}\}$, for some $\mathtt{d}\in\mathbb{N}$. Each particle is assigned a type $i\in\mathcal{I}$ and moves in $\mathbb{R}$ according to a type-dependent Lévy process. They randomly either branches, generating offspring that all inherit its current type, or switches to a new type. All particles evolve independently of one another. Specifically, the model is described as follows.
  • (Movement) For each particle of type $i\in\mathcal{I}$, its movement is governed by an $\mathbb{R}$-valued Lévy process $(\chi_{i}(t),t\geq 0)$. For $\theta\geq 0$, when the exponential moment $\mathbb{E}[e^{-\theta\chi_{i}(1)}]$ is finite, we define the Laplace exponent $\phi_{i}$ such that
@@
    $\mathcal{A}_{i}f(x)=\frac{\sigma_{i}^{2}}{2}\partial_{xx}f(x)+\mathtt{a}_{i}\partial_{x}f(x)+\int_{\mathbb{R}}\Big[f(x+y)-f(x)-y\mathbbm{1}_{\left\{|y|<1\right\}}\partial_{x}f(x)\Big]\,\Lambda_{i}(\mathrm{d}y).$ (1.3)
@@
    Note that, if $\beta_{j}=0$ for some $j\in\mathcal{I}$, then a particle cannot split at type $j$ and the offspring law is meaningless; so in this case we just fix by convention $\mu_{j}(1)=1$.
  • (Switching types) Each particle changes its type according to a continuous-time Markov chain $\Theta$ on $\mathcal{I}$, with intensity matrix $Q=(q_{ij})_{i,j\in\mathcal{I}}$, and possibly makes a jump at the time when it changes type. More precisely, at rate $q_{ij}\geq 0$, a particle of type $i$ switches to the type $j\in\mathcal{I}$, and at the same time makes a jump in space according to the law of a real-valued random variable $U_{ij}$, with convention that $U_{ij}=0$ if $q_{ij}=0$, and that $U_{ii}=0$. For $i\in\mathcal{I}$, let $q_{i}=-q_{ii}=\sum_{j\neq i}q_{ij}$. When $\mathbb{E}[e^{-\theta U_{ij}}]$ is finite, we define
@@
The theory of MAPs is well-established and has prominent applications in e.g. classical applied probabilistic models for queues. We refer to [4, Chapter XI], [30, Chapter 2] and [19, Appendix] for detailed discussions. The following proposition is well-known, giving a standard representation for MAPs.
@@ 1.3 Main results
Let $\theta\in[0,\bar{\theta})$. By the Perron-Frobenius (PF) theorem (see [53, Theorem 1.1] or [26, Theorem 8.3.4]), since $M^{(\theta)}(t)$ is a matrix with positive entries, the matrix $\mathcal{M}(\theta)$ has a PF eigenvalue $\lambda(\theta)$. This eigenvalue $\lambda(\theta)$ is real and larger than the real part of any other eigenvalues of $\mathcal{M}(\theta)$, and its associated right eigenvector $\vec{V}(\theta)=(V_{i}(\theta),i\in\mathcal{I})$ has strictly positive entries. Furthermore, because the entries of $\mathcal{M}(\theta)$ are infinitely differentiable on $(0,\bar{\theta})$, the PF eigenvalue $\lambda(\theta)$ and the corresponding eigenvector $\vec{V}(\theta)$ are also infinitely differentiable on this interval. The properties above of the PF eigenvalue are introduced in Lemma 2.3 and Lemma 2.4.
We observe that the number of particles $\left(\sum_{u\in\mathcal{N}_{t}}\mathbbm{1}_{\left\{J_{u}(t)=1\right\}},\ldots\sum_{u\in\mathcal{N}_{t}}\mathbbm{1}_{\left\{J_{u}(t)=\mathtt{d}\right\}}\right)_{t\geq 0}$, forms a continuous-time multitype branching process (see e.g. [5]): for each particle of type $i\in\mathcal{I}$, it is replaced by $k$ particles of type $i$ at rate $\beta_{i}\mu_{i}(k)$ for $k\geq 0$; whereas it is replaced by one type $j$ particle at rate $q_{ij}$ for $j\neq i$. When $\theta=0$, $\mathcal{M}(0)=Q+\mathrm{diag}(\beta_{1}m_{1}-\beta_{1},\ldots,\beta_{\mathtt{d}}m_{\mathtt{d}}-\beta_{\mathtt{d}})$ serves as the generator of the first moment semigroup of $\left(\sum_{u\in\mathcal{N}_{t}}\mathbbm{1}_{\left\{J_{u}(t)=1\right\}},\ldots\sum_{u\in\mathcal{N}_{t}}\mathbbm{1}_{\left\{J_{u}(t)=\mathtt{d}\right\}}\right)_{t\geq 0}$ and the PF eigenvalue $\lambda(0)$ determines the extinction behaviour with a phase transition. Denote the survival event by
@@
By [5, Theorem 2], the vector $\vec{\mathbf{q}}:=(\mathbf{q}_{1},\dots,\mathbf{q}_{\mathtt{d}})$ is the unique solution in $0,1]^{\mathtt{d}}\setminus\{(1,\dots,1)\}$ to the equation:
@@
The process also satisfies the “positive regularity” and “non-singularity” in [25, Theorem 2.7.1] and [25, Corollary 1 of Theorem 2.7.2]. Therefore the unique solution $\vec{\mathbf{q}}$ also lies in $0,1)^{\mathtt{d}}$. Whereas when $\lambda(0)\leq 0$, the extinction happens $\mathbf{P}_{x,i}$-a.s. $\forall x\in\mathbb{R},i\in\mathcal{I}$, i.e. $\mathbf{P}_{x,i}(\mathscr{S})=0$.
In what follows, we will always work in the supercritical case $\lambda(0)>0$. Summarizing, we assume that
@@
Theorem 1.4 (Additive martingale).
@@
By the main Theorem of [33], the function $q\mapsto\lambda(q)$ is strictly convex on $(0,\bar{\theta})$. It follows that $q\mapsto\frac{\lambda(q)}{q}$ either attains its unique minimum in the interval $(0,\bar{\theta})$ or approaches the infimum at the boundaries 0 or $\bar{\theta}$ (note that $\bar{\theta}$ could be $\infty$). The next assumption specifies that the minimum is attained in the interior.
@@
The main tool we use in the proof of Theorem 1.4 is the spine decomposition. The spine decomposition is a powerful technique in the study of branching processes. Typically, it contains two steps. The first step is a change of measure, and the second step is a deconstruction of the process in the new probability measure. For the details of the spine decomposition on classic branching systems, one can refer to [41, 54] for applications on branching random walks (BRWs), [37] for applications on branching Brownian motions (BBMs), and [50] for applications on branching Markov processes (BMPs). In our model, we established a spine decomposition, which is introduced formally in Section 2.3, and proved in Section 5. This framework provides a powerful tool for deeper analysis of the model, which we employ here to study the derivative martingale and the travelling waves.
@@
$Z_{\theta}(t):=\sum_{u\in\mathcal{N}_{t}}e^{-\theta X_{u}(t)-\lambda(\theta)t}\left[V_{J_{u}(t)}(\theta)(X_{u}(t)+\lambda^{\prime}(\theta)t)-V^{\prime}_{J_{u}(t)}(\theta)\right],\qquad t\geq 0.$ (1.13)
Note that we intuitively have $Z_{\theta}(t):=-\partial_{\theta}W_{\theta}(t)$. At criticality $\theta=\theta^{*}$, the derivative martingale is a crucial object for understanding the fine structure of the frontier in branching systems; see, for example, [1, 54] for applications on branching random walks, [37, 38] for applications on BBMs and [51] for applications on multitype BBMs.
@@
The FKPP equation is a well-known reaction-diffusion equation that appears in population genetics and ecology. Its connection to branching processes was first studied via probabilistic tools in [45]. Since then, it has been widely studied in probability, particularly in branching systems, see, for example, [13], [35], [37], [38] and [51]. In the study of FKPP equation, the existence, uniqueness and the asymptotic of the travelling wave solutions are particularly interested. One can find a “common pattern” in these works of FKPP equations. In the supercritical regime, the travelling wave solutions are often strongly connected to the limit of the additive martingale; while in the critical regime, it is related to the limit of the derivative martingale; while in the subcritical regime, there is no travelling wave solution.
@@
$\frac{\partial\mathbf{u}(t,x,i)}{\partial t}=$ $\,\mathcal{A}_{i}\mathbf{u}(t,x,i)+\sum_{j\neq i}q_{ij}\left(\int_{\mathbb{R}}\mathbf{u}(t,x\!+\!y,j)\mathbb{P}(U_{ij}\!\in\!\mathrm{d}y)-\mathbf{u}(t,x,i)\right)$ (1.15)
$\,+\beta_{i}\big(g_{i}(\mathbf{u}(t,x,i))-\mathbf{u}(t,x,i)\big).$ (1.16)
@@
$\mathbf{u}(t,x,i)=$ $\,\mathbb{E}_{x,i}\left[\mathbf{u}(0,\xi_{i}(t),i)\right]+\int_{0}^{t}\mathbb{E}_{x,i}\Bigg[\beta_{i}g_{i}\left(\mathbf{u}\left(s,\xi_{i}(t\!-\!s),i\right)\right)+$ (1.17)
$\sum_{j\neq i}q_{ij}\int_{\mathbb{R}}\mathbf{u}\left(s,y\!+\!\xi_{i}(t\!-\!s),j\right)\mathbb{P}(U_{ij}\in\mathrm{d}y)-(q_{i}+\beta_{i})\mathbf{u}\left(s,\xi_{i}(t\!-\!s),i\right)\Bigg]\mathrm{d}s.$ (1.18)
@@
$\mathcal{T}_{1}=\left\{f\;\middle|\;\forall i\in\mathcal{I},x\mapsto f(x,i)\text{ is non-decreasing in }x,\right$. (1.19)
$\hskip 60.00009pt\left.\lim_{x\to-\infty}f(x,i)=\mathbf{q}_{i}\text{ and }\lim_{x\to\infty}f(x,i)=1\right\}.$ (1.20)
@@
    (Subcritical regime) There are no travelling wave solutions with speed $\rho<\frac{\lambda(\theta^{*})}{\theta^{*}}$.